OpenStudy (anonymous):

what is the height of the cylinder. volume=271.4 in^3 , radius = 6 in

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

do \[\frac{271.4}{36\pi}\] probably need a calculator

OpenStudy (anonymous):

hey. I really need help with some math problems. I think the answer is 7.2?

OpenStudy (misty1212):

http://www.wolframalpha.com/input/?i=271.4%2F%2836pi%29 close to \(2.4\) i see

OpenStudy (misty1212):

no not \(7.2\)

OpenStudy (anonymous):

oh oops I did the math wrong

OpenStudy (anonymous):

find the volume of the cylinder in terms of pi. h=6in r=3in

OpenStudy (misty1212):

in terms of \(\pi\) lol

OpenStudy (dmndlife24):

Formula...V = (pi) r^2 h

OpenStudy (misty1212):

do \[6\times 9\pi\]

OpenStudy (misty1212):

"in terms of \(\pi\)" means just write \(54\pi\)

OpenStudy (anonymous):

okay but what is ^3

OpenStudy (misty1212):

it is \[V=\pi\times r^2\times h\] in your example \(h=6,r=3\) so \(r^2=3^2=9\)

OpenStudy (anonymous):

ohhh thanks. find the volume of the cylider in terms of pi. h=11m r=1.9m

OpenStudy (misty1212):

oh you mean what is the cubed doing there? the volume is a cubed unit, so \(54\pi\) cubed, whatever the units are

OpenStudy (misty1212):

same as the last one \[11\times 1.9^2\times \pi\]

OpenStudy (misty1212):

they probably want you to compute \(11\times 1.9^2\)

OpenStudy (misty1212):

http://www.wolframalpha.com/input/?i=11*1.9%5E2

OpenStudy (anonymous):

so is it 75.449?

OpenStudy (misty1212):

\[39.71\pi\] is what they are looking for

OpenStudy (anonymous):

oh once again I did the math wrong. find the volume of the given prism length=13.1cm width=4.7cm height=8.8cm

OpenStudy (misty1212):

no idea, sorry i don't really know any geometry you want me to google volume of a prism?

OpenStudy (anonymous):

yes any help is needed

OpenStudy (misty1212):

if it is a rectangular prism, aka a box, multiply \[13.1\times 4.7\times 8.8\]

OpenStudy (misty1212):

http://www.wolframalpha.com/input/?i=13.1*4.7*8.8

OpenStudy (anonymous):

541.8?

OpenStudy (misty1212):

looks like you "did the math" right this time

OpenStudy (anonymous):

thanks

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

what are the lateral area and the surface area of the cone. r=12ft l=21ft

OpenStudy (misty1212):

lateral area is \(\pi\times 12\times 21\)

OpenStudy (anonymous):

and surface area?

OpenStudy (misty1212):

googled it, got \[\pi r^2+\pi\times r\times l\]

OpenStudy (misty1212):

so if google is right \[\pi \times 12^2+\pi \times 12\times 21\]

OpenStudy (anonymous):

2,598.8?

OpenStudy (misty1212):

which i guess you could write as \[396\pi\]

OpenStudy (misty1212):

if you are supposed to answer "in terms of \(\pi\)

OpenStudy (anonymous):

this one isn't in terms of pi. so its 1,244 right?

OpenStudy (misty1212):

yeah

OpenStudy (anonymous):

thanks. can you help with 5 more?

OpenStudy (misty1212):

quickly maybe mayweather fight is coming on soon

OpenStudy (anonymous):

hahaha who r u cheering for?

OpenStudy (misty1212):

mayweather i guess he will win anyways i think

OpenStudy (misty1212):

lets see if we can do a couple more quick like

OpenStudy (anonymous):

find the SA of the cylinder r=12in l=17in

OpenStudy (anonymous):

this can be the last one

OpenStudy (misty1212):

\[2\pi rh +2\pi r^2\]

OpenStudy (misty1212):

\[2\pi \times 12\times 17+2\pi \times 144\]

OpenStudy (misty1212):

with or without \(\pi\)?

OpenStudy (anonymous):

without

OpenStudy (misty1212):

2186.5 looks like

OpenStudy (misty1212):

we can do one moreq

OpenStudy (anonymous):

thanks okay

OpenStudy (anonymous):

what are the lateral area and surface area of the box to the nearest whole number? use the large 17cm x 5 cm rectangles on the top and bottom of the bases. l= 17 w= 5 h=3.12

OpenStudy (misty1212):

lateral area \[(2l+2w)h\]

OpenStudy (misty1212):

\[(2\times 17+2\times 5)\times 3.12\]

OpenStudy (misty1212):

http://www.wolframalpha.com/input/?i=%282*17%2B10%293.2 141 to the nearest whole number

OpenStudy (misty1212):

surface area \[2(17\times 5+5\times 3.2+3.2\times 17)

OpenStudy (misty1212):

\[2(17\times 5+5\times 3.2+3.2\times 17)\]

OpenStudy (anonymous):

thanks

OpenStudy (misty1212):

yw good luck \[\color\magenta\heartsuit\]